Question  9

The marginal cost function for a good is given by

MC = −Q2 + 80Q.

Find the total cost function if fixed costs are   500.


  1. -Q3/3 +40 + 50
  2. Q3/3 +40Q2 + C
  3. -Q3/3 +80Q2 + C
  4. Q3 +80Q2 + 500
1
Expert's answer
2021-07-05T18:12:01-0400

"MC = -Q^2+80Q\\
\int MC dQ = \int (-Q^2 +80Q)dQ = -\cfrac{Q^3}{3}+40Q^2+C"

So total cost function if cost are 500 is:

"TC = -\cfrac{Q^3}{3}+40Q^2+500"